# 3D variational autoencoder for fingerprinting microstructure volume elements

### v2

@@ 3.3.1 Linear Paths
$z_{n}=z_{1}+\frac{n+1}{N}z_{2},\hskip 10.00002ptn=1,\dots,N.$
The $z_{n}$ can be passed through the trained decoder to reconstruct the microstructures, $\hat{x}$, along the path. Figure 7 shows an example of such a path with $N=8$. This shows a smooth transition between each microstructure along the path, visually confirming the continuity of the latent space. Continuity in the latent space is a desirable property for downstream tasks, such as parameterising effective VE properties by their fingerprint, since it guarantees smooth property mappings.
@@ 3.3.2 Dimensionality Reduction
Principal component analysis (PCA) [49] was used to reduce the fingerprints down to two dimensions. This is done for visualisation purposes only. The explained variance was shown to only be fully captured if the first 200 principal components were used, so this 2D representation is by no means the complete picture. It does, however, provide some insight into which microstructural information is differentiated by the resulting fingerprints. Spectral clustering is used to verify these findings in Section 3.3.3.
@@
The correlation between texture and the latent space can be seen in Figures 8(c) and 8(d), which show fingerprints from VEs with some common textures and VEs with some specific normalised Euler angles, chosen such that they are sufficiently far away from zero. As discussed in Section 2.1.2, whilst generating these VEs, the orientation is forced to be strictly positive, and the absolute value is taken of any noise added to an Euler angle close to zero. This avoids cases where the resulting angle would wrap around from 0 to $2\pi$, but results in an artificially smaller variance in cases where the Euler angle is close to zero. This can be seen in Figure 8(c), where the Cube and Goss textures show much tighter clustering than the other cases. The set of edge cases with normalised Euler angles further away from zero, with a relatively small variance, were also analysed in the same way to show more distinct separation and tighter clustering. This can be seen in Figure 8(d).
@@ 3.4 Surrogate Model Results
Figure 9(b) shows the stress response for four sample VEs picked randomly from the validation dataset. These were selected to demonstrate that VEs with a stress response around the mean are predicted with a high accuracy, whereas stress responses nearer the extreme values suffer from worse prediction accuracy. However, these edges are still captured somewhat, showing that there is indeed some microstructural dependence being afforded by the VAE fingerprints. The discrepancy in accuracy for these edge cases is potentially due to a lack of representation in the training dataset. It is also worth noting here that the stress response predictions are almost identical for all VEs for the first five strain increments. However, these generally fall within the elastic region of the material, in which there is less of an impact from the anisotropy. Figure 9(c) shows the mean relative error, i.e. $\mathbb{E}(|y-\hat{y}|)$, where $y$ denotes the ground truth stress and $\hat{y}$ is the corresponding prediction, with the standard deviation shown as a shaded region on the plot. The mean relative error across all strain increments is 2.75 $\pm$ 0.46 MPa, and is fairly constant across all strain increments except for a drop in error within the elastic region, where anisotropy has less of an effect.
@@ 4 Discussion
We have shown that a VAE based on ResNet18, modified with 3D convolutions, can be trained to encode and accurately reconstruct microstructure VEs with an Euler angle triplet at each voxel. However, a relatively large dataset was required for training and care must be taken to account for the symmetries involved. To ensure that the training process is adaptable to all crystal types, symmetry was removed as a preprocessing step. It is possible to construct a loss function that takes the symmetries into account, but the Euler angle fundamental zones have a non-trivial geometry, so it is usual to project the Euler angles to quaternions before mapping to the fundamental zone, and this is an expensive operation to perform repeatedly throughout the training process. Furthermore, dealing with symmetries in the loss function can result in a discontinuous loss, which should be avoided for backpropagation. Dealing with symmetries as a preprocessing step allows for the loss to be continuous, but it is important to note that a VAE trained on a dataset with a specific fundamental zone, say cubic or hexagonal, is restricted to operating on unseen data constructed from the same fundamental zone. If symmetry is not considered at all, the VAE can be trained to operate on any crystal structures, making it more generalisable. However, the latent space then views symmetric variants of an Euler angle as completely different angles and, as such, textured VEs may not cluster in the latent space as one might expect. Additionally, this would result in learning a significantly more complex mapping, requiring greater optimisation effort as well as larger model parameterisation. Euler angles were used primarily as they allow for standard convolutions to be applied to each angle in the triplet independently, as separate image channels. They are also simple to normalise and any reconstructions output from the VAE would be valid and applicable to being input into the loss function. If, say, quaternions were used instead, the reconstructions would not necessarily yield valid quaternions without some additional constraint in the loss function to force the $L_{2}$ norm to equal one.
@@
Some key parameters related to the VAE architecture (latent space dimensionality and ResNet depth) were optimised with a small grid search, resulting in a dimensionality reduction of over 3000×3000$\times$3000 × and a mean misorientation between inputs and reconstructions of $0.009\pm 0.006$ for unseen VEs that are statistically equivalent to the training data. This grid search was performed over more thorough optimisation algorithms due to the computational cost of training the VAE with the resources available. Once trained, the VAE was tested on some edge cases. These included grain morphologies and textured VEs that were not included in the training dataset. For the morphology cases, it was shown that the VAE generalises well to VEs with larger grains than those sampled to construct the training data and, also, to elongated grains as opposed to equiaxed grains, with mean misorientations of $0.006\pm 0.007$ and $0.009\pm 0.009$, respectively. However, in cases where the grain size was much smaller than in the training dataset, the reconstructions were very poor, with a mean misorientation of $0.072\pm 0.107$. This could be due to the minimum resolution afforded by the convolution operations in the VAE, although it is more likely that this is an issue with extrapolating from the grain size in the training data, in this case. It is very possible that the larger and elongated grains are being interpreted as equiaxed grains, with similar morphologies to the training data and adjacent grains containing matching Euler angles.
@@
When keeping the grain morphology consistent with the training data and considering textured VEs as edge cases, the VAE generalises well in all cases with mean misorientations on the order of $10^{-3}$ or $10^{-4}$. Despite these textures not explicitly existing in the training data, the orientations sampled to construct them do exist in the training data, and so this is more reliant on interpolation of the latent space, which works well. Therefore, to construct a more general latent space for microstructure, sampling orientations uniformly across the Euler space leads to good generalisation, but higher variance on the grain size distribution across the training data is advisable, although this will likely result in the requirement for a larger training dataset.
@@
The trained VAE was used to construct a new training dataset that consisted of a set of fingerprints with associated stress responses under uniaxial tension, for a set of magnesium VEs. A fully connected network was trained as a microstructure-dependent surrogate for the CP simulations. This was shown to be somewhat effective, with the spread in stress response across the dataset being captured, although, the model would benefit from a larger training dataset and/or a more carefully curated dataset to yield a more uniform sampling from the stress space. Active learning could be employed to this effect, but it is difficult to identify microstructures that would result in edge cases a priori. However, given a large enough dataset, it could be sampled to remove bias towards the mean stress response. Once trained, the stress response can be predicted in $O(10^{-5})$ seconds, yielding a speed-up of $O(10^{6})$ over the CP simulation, with a mean relative error of 2.75 $\pm$ 0.46 MPa. This speed-up would become even more prominent if finer stress increments were used, although the course strain increments used here were deemed favourable in this case to enable a relatively large dataset to be constructed for training. It is important to note that, although we have established microstructure dependence here, we have neglected any dependence on the load path. Combining methods for fingerprinting VEs with a surrogate for predicting stress response from a given load path remains an avenue for further work. This could be achieved by training an RNN that takes a VAE fingerprint as input, to encode the initial microstructure configuration into the hidden state of the RNN, alongside inputting the strain tensor at each recurrent block. This could be trained to predict the stress response for a given microstructure VE and prescribed strain path. The speed-up that this would afford, compared to running a full CP simulation, would enable microstructure-dependent stress responses to be embedded into larger scale finite element models, at each node individually, bridging the gap between crystal plasticity and the component scale models.

### v3

@@ 3.3.1 Linear Paths
$z_{n}=z_{1}+\frac{n+1}{N}z_{2},\hskip 10.00002ptn=1,\dots,N.$ (6)
The $z_{n}$ can be passed through the trained decoder to reconstruct the microstructures, $\hat{x}$, along the path. Figure 9 shows an example of such a path between two volume elements in the test dataset that were not seen during training, with $N=8$. The images shown are slices through the centre of each 3D reconstruction and show a smooth transition between each microstructure along the path, visually confirming the continuity of the latent space. Continuity in the latent space is a desirable property for downstream tasks, such as parameterising effective VE properties by their fingerprint, since it guarantees smooth property mappings. In some cases along the path shown in Figure 9, there appears to be some discontinuities with individual grains suddenly switching orientation. This is due to the step size taken between reconstructions in the latent space (smaller steps yields smoother transformations). To verify this, we have plotted a path with a finer step size for one of these cases, which shows that the change in orientation exhibits smooth transition by engulfing the original grain, rather than there being a sudden rotation in the orientation space.
@@ 3.3.2 Dimensionality Reduction
PCA [50] was used to reduce the fingerprints down to two dimensions for visualisation purposes only and not to further compress the fingerprints for downstream tasks. The explained variance was shown to only be fully captured if the first 200 principal components were used, so this 2D representation is by no means the complete picture. It does, however, provide some insight into which microstructural information is differentiated by the resulting fingerprints. Spectral clustering is used to verify these findings in Section 3.3.3.
@@
The correlation between texture and the latent space can be seen in Figures 10(c) and 10(d), which show fingerprints from VEs with some common textures and VEs with some specific normalised Euler angles, chosen such that they are sufficiently far away from zero. As discussed in Section 2.1.2, whilst generating these VEs, the orientation is forced to be strictly positive, and the absolute value is taken of any noise added to an Euler angle close to zero. This avoids cases where the resulting angle would wrap around from 0 to $2\pi$, but results in an artificially smaller variance in cases where the Euler angle is close to zero. This can be seen in Figure 10(c), where the Cube and Goss textures show much tighter clustering than the other cases. The set of edge cases with normalised Euler angles further away from zero, with a relatively small variance, were also analysed in the same way to show more distinct separation and tighter clustering. This can be seen in Figure 10(d).
@@ 3.4 Surrogate Model Results
Figure 11(b) shows the stress response for four sample VEs picked randomly from the validation dataset. These were selected to demonstrate that VEs with a stress response around the mean are predicted with a high accuracy, whereas stress responses nearer the extreme values suffer from worse prediction accuracy. However, these edges are still captured somewhat, showing that there is indeed some microstructural dependence being afforded by the VAE fingerprints. The discrepancy in accuracy for these edge cases is potentially due to a lack of representation in the training dataset. It is also worth noting here that the stress response predictions are almost identical for all VEs for the first five strain increments. However, these generally fall within the elastic region of the material, in which there is less of an impact from the anisotropy. Figure 11(c) shows the mean relative error, i.e. $\mathbb{E}(|y-\hat{y}|)$, where $y$ denotes the ground truth stress and $\hat{y}$ is the corresponding prediction, with the standard deviation shown as a shaded region on the plot. The mean relative error across all strain increments is 2.75 $\pm$ 0.46 MPa, and is fairly constant across all strain increments except for a drop in error within the elastic region, where anisotropy has less of an effect. Note that the standard deviation reported here is the standard deviation of the mean relative error across each time step and not the mean standard deviation of the relative error across each time step, which is 2.35.
@@ 4 Discussion
We have shown that a VAE based on ResNet18, modified with 3D convolutions, can be trained to parameterise microstructure for input into a downstream microstructure-dependent surrogate model for predicting material behaviour - in this case, a simple uniaxial CP simulation with fixed loading. It is foreseen that the pipeline presented here could be adapted by switching out the predictive surrogate with any model of interest, without the need to retrain a microstructure encoder, as one would need to do for a full end-to-end model. The VAE was shown to accurately reconstruct microstructure VEs with an Euler angle triplet at each voxel, however, a relatively large dataset was required for VAE training and care must be taken to account for the symmetries involved. The reconstructions are used to train the encoder in a self-supervised manner, which can be used to parameterise microstructure in a predictive model, and the trained decoder is not used in any downstream tasks, other than to provide some visual aid towards understanding the latent space that has been constructed. To ensure that the training process is adaptable to all crystal types, symmetry was removed as a preprocessing step. It is possible to construct a loss function that takes the symmetries into account, but the Euler angle fundamental zones have a non-trivial geometry, so it is usual to project the Euler angles to quaternions before mapping to the fundamental zone, and this is an expensive operation to perform repeatedly throughout the training process. Furthermore, dealing with symmetries in the loss function can result in a discontinuous loss, which should be avoided for backpropagation. Dealing with symmetries as a preprocessing step allows for the loss to be continuous, but it is important to note that a VAE trained on a dataset with a specific fundamental zone, say cubic or hexagonal, is restricted to operating on unseen data constructed from the same fundamental zone. If symmetry is not considered at all, the VAE can be trained to operate on any crystal structures, making it more generalisable. However, the latent space then views symmetric variants of an Euler angle as completely different angles and, as such, textured VEs may not cluster in the latent space as one might expect. Additionally, this would result in learning a significantly more complex mapping, requiring greater optimisation effort as well as larger model parameterisation. Euler angles were used primarily as they allow for standard convolutions to be applied to each angle in the triplet independently, as separate image channels. They are also simple to normalise and any reconstructions output from the VAE would be valid and applicable to being input into the loss function. If, say, quaternions were used instead, the reconstructions would not necessarily yield valid quaternions without some additional constraint in the loss function to force the $L_{2}$ norm to equal one.
@@
Some key parameters related to the VAE architecture (latent space dimensionality and ResNet depth) were optimised with a small grid search, resulting in a dimensionality reduction of over 3000×3000$\times$3000 × and a mean misorientation between inputs and reconstructions of $0.031\pm 0.004$ for unseen VEs that are statistically equivalent to the training data. This grid search was performed over more thorough optimisation algorithms due to the computational cost of training the VAE with the resources available. Once trained, the VAE was tested on some edge cases. These included grain morphologies and textured VEs that were not included in the training dataset. For the morphology cases, it was shown that the VAE generalises well to VEs with larger grains than those sampled to construct the training data and, also, to elongated grains as opposed to equiaxed grains, with mean misorientations of $0.033\pm 0.010$ and $0.036\pm 0.009$, respectively, which is comparable to the reconstruction accuracy on the training data. However, in cases where the grain size was much smaller than in the training dataset, the reconstructions were very poor, with a mean misorientation of $0.063\pm 0.001$. This is due to the larger and elongated grains effectively being a subset of the training data, which can be constructed by sampling grains that are statistically equivalent the training data, but with adjacent grains sharing the same crystallographic orientation. Whereas the smaller grains are only achievable through extrapolation away from the training dataset. Therefore, the VAE generalises well to unseen within-distribution samples, but not outside the distribution that it is trained on. The VAE was retrained on a dataset containing smaller grains with two distinct mean ESDs to add more variability to the dataset and this showed that suitable representation in the training data and increasing the size of the latent space can lead to better encoding and reconstruction of the small grain edge cases.
@@
When keeping the grain morphology consistent with the training data and considering textured VEs as edge cases, the VAE generalises well in all cases with mean misorientations comparable to those for the training data. Despite these textures not explicitly existing in the training data, the orientations sampled to construct them do exist in the training data, and so this is more reliant on interpolation of the latent space, which works well. Therefore, to construct a more general latent space for microstructure, sampling orientations uniformly across the Euler space leads to good generalisation, but higher variance on the grain size distribution across the training data is advisable, although this will likely result in the requirement for a larger training dataset.
@@
Once hyperparameters have been selected, the heavyweight training of the VAE is only required to be performed once and allows for a more lightweight surrogate model to be appended, which can be swapped out depending on the task. This is much more efficient than training a new end-to-end model each time a new predictive model is of interest. In this case, the trained VAE was used in conjunction with DAMASK to construct a new training dataset that consisted of a set of fingerprints with associated volume-averaged stress responses under uniaxial tension, for a set of magnesium VEs, as a proof of concept for applying VAE fingerprints to produce a microstructure-dependent surrogate model. A fully connected network was trained as a microstructure-dependent surrogate for the CP simulations. This was shown to be somewhat effective, with the spread in stress response across the dataset being captured, although, the model would benefit from a larger training dataset and/or a more carefully curated dataset to yield a more uniform sampling from the stress space. Active learning could be employed to this effect, but it is difficult to identify microstructures that would result in edge cases a priori. However, given a large enough dataset, it could be sampled to remove bias towards the mean stress response. Once trained, the stress response can be predicted in $O(10^{-5})$ seconds, yielding a speed-up of $O(10^{8})$ over the CP simulation, with a mean relative error of 2.75 $\pm$ 0.46 MPa. This speed-up would become even more prominent if finer stress increments were used, although the course strain increments used here were deemed favourable in this case to enable a relatively large dataset to be constructed for training. It is important to note that, although we have established microstructure dependence here, we have neglected any dependence on the load path and developing a surrogate for predicting stress response from a given load path remains an avenue for further work. However, the advantage of training a VAE separately to the predictive model, rather than a fully end-to-end model, is that the trained VAE encoder is application agnostic and can be bolted on to any feasible microstructure-dependent surrogate model as a preprocessing step. This could be achieved by training an RNN that takes a VAE fingerprint as input, to encode the initial microstructure configuration into the hidden state of the RNN, alongside inputting the strain tensor at each recurrent block. This could be trained to predict the stress response for a given microstructure VE and prescribed strain path. The speed-up that this would afford, compared to running a full CP simulation, would enable microstructure-dependent stress responses to be embedded into larger scale finite element models, at each node individually, bridging the gap between crystal plasticity and the component scale models.
@@ 5 Conclusions
  • The trained VAE can be used as a preprocessing step for parameterising microstructure for microstructure-dependent surrogate models.